# Thompson Sampling via Fine-Tuning of LLMs

### v1

@@ 5.1 Baselines
We compare against four baselines. For fair comparison, the reward model (if present) is always a Gaussian process over the same (deep) embeddings. As in Li et al. (2022), Unguided Generation samples directly from the pre-trained LLM without feedback, serving as a non-adaptive baseline. As in Kristiadi et al. (2024) and Ranković & Schwaller (2025), Post-Generation TS performs Thompson sampling over a fixed subset of candidates (of size 1000), here generated by the LLM prior to optimization. While computationally efficient, PG-TS is constrained by its static candidate pool and, as our experiments show, consistently underperforms ToSFiT across all tasks. Finally, Actor Critic (Barto et al., 1983) maximizes $\mathbb{E}_{x\sim\pi^{\theta}}[\mu_{x}]$ rather than $\mathcal{V}$, and Soft Actor Critic (Haarnoja et al., 2018) additionally adds entropy regularization, i.e., maximizes $\mathbb{E}_{x\sim\pi^{\theta}}[\mu_{x}-\alpha\ln\pi^{\theta}_{x}]$. ToSFiT is more performant than either, since entropy regularization is only applied toward unexplored regions where $\sigma_{x}$ is sufficiently large.
@@
Figure 4: Unguided Generation quickly plateaus, and classical Bayesian optimization over a fixed LLM-generated candidate set saturates early (Post-Generation TS). In contrast, ToSFiT explores the full solution space and consistently discovers higher-reward candidates. Via optimism in the face of uncertainty, it explores more efficiently than Actor Critic and Soft Actor Critic.
5.2 Settings
FAQ Refinement
is a natural language task that tests the algorithm’s ability to optimize text based on semantic alignment. We ask a Qwen3-1.7B model (Yang et al., 2025) to write a frequently-asked-questions (FAQ) response. The reward is modeled as the alignment to an unknown ground-truth response, judged by the Qwen3-Embedding-0.6B model (Zhang et al., 2025). As a deep kernel, we adopt a linear kernel over the first 256 entries of the embeddings of Qwen3-Embedding-0.6B. The search space consists of all token sequences, which is exponentially large in the response length.
Protein Search
explores the challenge of designing thermally stable proteins—a task with significant implications for drug development and industrial biotechnology. The goal is to identify amino acid sequences that exhibit high thermal stability, a property that enhances protein robustness and shelf life. Note that with 20 standard amino acids in the human body and sequence lengths of $100$ and above, the search space exceeds the number of atoms in the observable universe. We sample amino acid sequences from ProtGPT2 (Ferruz et al., 2022) and score them according to their negative thermal instability index (Guruprasad et al., 1990). To predict the thermal stability and assess epistemic uncertainty, we adopt a Gaussian process with linear kernel over the mean token embeddings from ProtGPT2 projected to the unit sphere.
Quantum Circuit Design
is the task of designing quantum circuits that prepare low-energy quantum states in unknown environments. The challenge lies in navigating a vast, discrete space of valid quantum programs, where entanglement and gate structure critically influence performance. To generate Qiskit circuits (Javadi-Abhari et al., 2024), we use a Qwen2.5-Coder-1.5B model (Hui et al., 2024) prompted to fill-in-the-middle (Bavarian et al., 2022) after initializing six disentangled qubits in the zero state. As reward, we consider the negative energy of the prepared state under an unknown Hamiltonian with strong interaction terms, requiring entanglement for optimal performance. The feature map for the linear Gaussian process over rewards consists of a code validity bit as well as all two-qubit Pauli observables, which can be efficiently simulated using quantum computers.
Experimental Hyperparameters
@@ 5.2 Settings
ToSFiT requires minimal task-specific hyperparameter tuning. Across all experiments, a burn-in period of $m=16$ steps is adopted and an exploration bonus of $4.0$ is applied to the prior amplitude inferred via MLM. The number of steps of gradient ascent per step of Bayesian optimization, $c$, is fixed at $1$ to minimize computational overhead compared to Unguided Generation. To ensure stable estimation of VBOS gradients, the generation batch size $B$ is always set to $16$. We conduct unbatched Bayesian optimization ($b=1$) unless otherwise specified. Learning rates are the only hyperparameters tuned per task: $1.0\times 10^{-7}$ for FAQ refinement, $1.0\times 10^{-5}$ for Protein Search, and $5.0\times 10^{-6}$ for Quantum Circuit Design. The baseline Soft Actor Critic requires an additional entropy regularization coefficient as hyperparameter and is swept across $\{0.01,0.1,1.0,10.0\}$ for each experiment. We report on the best configuration, i.e., $0.1$ for FAQ Refinement, $1.0$ for Protein Search, and $10.0$ for Quantum Circuit Design.
@@
To reduce the memory overhead of ToSFiT compared to Unguided Generation, we do not adopt parameter-wise optimization statistics (Kingma & Ba, 2015). Instead, we use vanilla stochastic gradient descent with a fixed learning rate and without momentum. We do not store the gradients, but instead apply them directly during the backward pass. Also, while we perform fully-batched auto-regressive generation, the forward & backward passes to compute $\tfrac{d}{d\theta}\ln\pi_{\theta}(x_{i})$ are performed in unbatched fashion with checkpointing. This crucially brings down memory consumption with little impact on runtime. All experiments were run on a single NVIDIA A100 GPU with 80 Gigabytes of VRAM. Section C.1 analzes the latency overhead of ToSFiT compared to Unguided Generation, and only reports a 19% overhead.
5.3 ToSFiT Overcomes the Reward Plateau (A)
Consider Figure 4. In all three tasks—FAQ Response Refinement, Protein Search, and Quantum Circuit Design—the best-seen reward of unguided generation quickly saturates at a suboptimal level. Classical Bayesian optimization over a fixed subset of generated candidates identifies a good solution more efficiently, but remains confined to the initial sample pool and cannot escape the performance plateau. In contrast, ToSFiT performs Bayesian optimization over the full solution space, enabling it to discover significantly better candidates beyond the reach of static sampling strategies.
5.4 Strong Initial Policy Accelerates Optimization (B)
Theorem 1 highlights the importance of initializing the sampling policy with strong prior knowledge. Figure 5 provides empirical support: masking the number of qubits in the prompt (ToSFiT weak context) leads to a significantly worse reward than using the full prior context (ToSFiT). Note that uniform initialization fails to produce valid code.
The theorem further suggests cautious adaptation to avoid drifting too far from the initialization. As shown in Figure 6, a large learning rate may initially improve performance by closely following VBOS, but eventually leads to forgetting the prior and stagnating as optimization becomes harder.
5.5 ToSFiT for Batched Bayesian Optimization (C)
6 Related Work
As a remedy for the intractable maximization of acquisition functions over combinatorially large discrete domains, Bal et al. (2025) assume a factorization into the cartesian product of much smaller domains. They then identify local maxima as configurations where changing any of the factors individually decreases the acquisition function, and query the black-box function with the highest scoring game equilibrium. In contrast to ToSFiT, the domain must be structured to admit a factorization into cartesian products, and their optimization strategy does not consider the strong priors available from pretrained large language models.
Sutter et al. (2025) train an in-context Thompson sampler (FIBO) that parameterizes the probability of maximality given a history of in-context observations. At inference, FIBO does not require any gradient updates and instead relies on in-context conditioning. However, their generator must be pre-trained to conduct Thompson sampling and thus cannot directly leverage large-scale generative pre-training. Moreover, their method is only demonstrated to work on synthetic function maximization, where classical Bayesian optimization strategies such as GP-UCB (Srinivas et al., 2010) and Thompson sampling (Russo et al., 2018) excel.
There is extensive prior work on relaxing Bayesian optimization over discrete space to continuous space using variational autoencoders (Kusner et al., 2017; Lu et al., 2018; Griffiths & Hernández-Lobato, 2020; Notin et al., 2021; Lee et al., 2025). Acquisition function maximization is then tractably performed via gradient ascent in the continuous latent space. However, these methods require a task-specific pre-training phase for the autoencoder, whereas ToSFiT can be applied zero-shot through prompt-conditioning, thus benefiting from large-scale industrial LLM pretraining.
Finally, Ranković & Schwaller (2025) learn deep kernel features for Bayesian optimization with Gaussian processes. By treating the neural feature map as a hyperparameter, they use gradient ascent to maximize the marginal likelihood of observations during Bayesian optimization in a fully online fashion. Experimentally, they find much better optimization trajectories compared to fixed feature maps. That said, they do not address the intractability of acquisition function maximization. Their method is complementary to ToSFiT and could be integrated to learn task-specific embeddings.
7 Conclusion
We demonstrate that Thompson sampling can be efficiently scaled to large unstructured discrete domains by parameterizing the probability of maximality with a generative policy. Our theoretical analysis reveals that VBOS already leverages the kernel structure of Gaussian processes and achieves the strong regret bound of standard Thompson sampling and GP-UCB. Extending these results to approximate VBOS highlights the importance of initializing the sampling policy according to strong priors from pre-training and adapting it cautiously to maintain prior knowledge. These insights are supported by empirical results across diverse tasks, where ToSFiT consistently outperforms the baselines, overcoming reward plateaus and naturally extending to batched optimization. Together, our findings demonstrate the potential of combining foundation models with principled Bayesian optimization to tackle complex, discrete search problems.
7.1 Limitations and Future Work
To ensure controlled comparisons, we evaluate ToSFiT under fixed feature maps—either derived from pre-trained embeddings or manually designed. A promising direction for future work is to learn deep, task-adaptive embeddings jointly with the Gaussian process, as in Ranković & Schwaller (2025), or replace the GP with more expressive reward models such as Bayesian neural networks or ensembles (Lakshminarayanan et al., 2017). Moreover, to reduce the computational and memory overhead introduced by fine-tuning, one could restrict updates to the last few layers of the generator. Finally, alternative strategies for optimizing the VBOS objective $\mathcal{V}(\pi)$—such as in-context conditioning—may offer a lightweight alternative to weight adaptation.
Ethics Statement
Reproducibility Statement
Apart from the code provided in the supplementary materials, pseudocode for ToSFiT is given in Algorithm 1. The baselines are also implemented in the code in a self-contained manner. The experimental environments for the three tasks considered is also provided in the code. Moreover, the provided README.md gives a step-by-step tutorial to reproduce all results in the paper. The hardware and software for running the experiments is described in detail in Section D of the Appendix. Furthermore, Section D.6 details the hyperparameter sweep conducted during development of this paper. In the evaluation, statistical significance is indicated via mean and standard error under $25$ seeds, set to $0,\ldots,24$. All theoretical statements are self-contained and include the complete set of assumptions. Both detailed proofs as well as proof sketches for the main results are provided in Section E of the Appendix. We distinguish between novel claims and prior statements by citing the appropriate works.
8 Acknowledgment
Appendix
Appendix A Standardized RLOO Is GRPO
Proposition 3.
Appendix B Constant-Time GP Inference
To ensure efficient runtime in large-scale settings and to integrate with deep neural features $\phi:X\to\mathbb{R}^{d}$, we consider Gaussian processes (GPs) with linear kernels over $\mathbb{R}^{d}$ (Rasmussen & Williams, 2006). As we explain in the following, our implementation requires a constant $\Theta(d^{2})$ memory and compute for conditioning on an observation, inferring posterior mean & variance at a point, and even conducting marginal likelihood maximization. In contrast to a straightforward implementation, it does not increase with the number of past observations.
B.1 Setting
B.2 Closed-Form MLM
Proposition 4.
B.3 Bypassing the Covariance Matrix
Appendix C Additional Experiments
C.1 Computational Cost of ToSFiT
Bayesian optimization is designed to trade the number of evaluations of an expensive reward function against the computationally demanding computation of the reward posterior. In this spirit, we emphasize sample efficiency as the main consideration. Nevertheless, as measured in wall-clock time in Table 1, the computational efficiency of our method can be made comparable to that of Unguided Generation.
Generation ToSFiT GP Update
$1.08\pm 0.05$ s $1.29\pm 0.05$ s $0.33\pm 0.06$ ms
Table 1: Wall-clock time per step of fully-batched ToSFiT ($b=B$) on an A100 GPU for Protein Search. The majority of the runtime is spent on candidate generation.
Note that in the fully-batched setting all generations are both used to estimate VBOS gradients and to provide observational feedback to the reward model. The 19% overhead observed in Table 1 stems from the additional non-auto-regressive backward pass.
Appendix D Experimental Details
To ensure reproducibility, we provide the setup for each of the experiments in Section 5. All experiments were conducted on compute nodes with an AMD EPYC 7763 64-Core CPU, 2TB RAM, and an NVIDIA A100-SXM4 GPU (80GB), running on Red Hat Enterprise Linux 9.4 with CUDA 12.4. Our Pytorch-based code (Paszke et al., 2019) is provided in Supplementary Materials, including an environment file specifying the necessary Python packages and their versions.
D.1 Baselines
For Unguided Generation, we simply generate and evaluate candidates according to the hyperparameters detailed in Section D.2. For Post-Generation TS, we first generate a fixed candidate pool using the large language model with the same hyperparameters, where the pool size is set to the number of steps of Bayesian optimization. Then, we conduct standard Thompson sampling on the candidate subset using a Gaussian process surrogate model derived from the (deep) task-dependent embeddings. For ToSFiT, we generate and evaluate candidates according to the same hyperparameters, but additionally follow Algorithm 1 to perform weight updates for consistency between the policy and the surrogate reward model, the same Gaussian process as in Post-Generation TS. The Actor Critic baseline follows ToSFiT, but maximizes $\mathbb{E}_{x\sim\pi^{\theta}}[\mu_{x}]$ rather than $\mathcal{V}$. Finally, the baseline Soft Actor Critic additionally adds entropy regularization to Actor Critic, i.e., maximizes $\mathbb{E}_{x\sim\pi^{\theta}}[\mu_{x}-\alpha\ln\pi^{\theta}_{x}]$. The hyperparameter $\alpha$ is swept for each experimental setting.
D.2 Experimental Details Shared Across Experiments
We run Bayesian optimization for $1000$ steps with a burn-in period of $16$ steps during which marginal likelihood maximization is conducted, but no gradients are applied (for ToSFiT) and queries are selected uniformly at random among the sub-selection (for Post-Generation TS). Marginal likelihood maximization is continuously applied during the full optimization trajectory, and an exploration bonus of $4.0$ is adopted. The noise-to-amplitude ratio $\sigma_{nar}$ is always set to $0.01$, reflecting the absence of noise in the measurements, but ensuring stable inversion of the observation covariance matrix. We sample from the language model with batch size $B=16$ and temperature set to $1.0$ while limiting the number of generated tokens to $512$. The number of fine-tune steps per batch of Bayesian optimization $c$ is universally set to $1$. Unless otherwise specified, unbatched Bayesian optimization is conducted. To optimize against the moving target provided by VBOS, vanilla stochastic gradient descent with neither momentum nor regularization is adopted. All experiments are repeated for $25$ random seeds.
D.3 FAQ Refinement

### v3

@@ 5.2 Baselines
We compare against seven baselines. For fair comparison, the reward model (if present) is always a Gaussian process over the same (deep) embeddings. As in Li et al. (2022), Unguided Generation samples directly from the pre-trained LLM without feedback, serving as a non-adaptive baseline. As in Kristiadi et al. (2024) and Ranković and Schwaller (2025), Post-Generation TS performs Thompson sampling over a fixed subset of candidates (of size 1000), here generated by the LLM prior to optimization. Actor Critic (Barto et al., 1983) maximizes $\mathbb{E}_{x\sim\pi^{\theta}}[\mu_{x}]$ rather than $\mathcal{V}$, and Soft Actor Critic (Haarnoja et al., 2018) additionally adds entropy regularization, i.e., maximizes $\mathbb{E}_{x\sim\pi^{\theta}}[\mu_{x}-\alpha\ln\pi^{\theta}_{x}]$. Evolutionary Search (Character) (Holland, 1975) treats solutions as sequences of text characters and applies biology inspired crossover and mutation operators. Evolutionary Search (LLM) (Romera-Paredes et al., 2024) prompts a language model to conduct crossover and mutation of solutions. Finally, FIBO (Sutter et al., 2025) performs Thompson sampling fully in-context of an LLM by attending to a list of evaluated candidate solutions.
5.3 ToSFiT Obtains State-of-the-Art Sample Efficiency
Figure 5: ToSFiT balances exploration with exploitation leading to stable policy diversity. Shown on Quantum Circuit Design, for other settings, see Section A.2.
Consider Figure 4. In all three tasks—FAQ Response Refinement, Protein Search, and Quantum Circuit Design—ToSFiT obtains state-of-the-art sample efficiency. Indeed, the best-seen reward of Unguided Generation quickly saturates at a suboptimal level. Classical Bayesian optimization over a fixed subset of generated candidates (Post-Generation TS) identifies a good solution more efficiently, but remains confined to the initial sample pool. Evolutionary Search as well as Actor Critics conduct undirected exploration without optimism in the face of uncertainty, i.e., an exploration bias guided by the reward potential of unexplored regions. As confirmed in Figure 5, the optimism of ToSFiT results in a more stable exploration-exploitation tradeoff than that of Actor Critics. FIBO is closest to ToSFiT in spirit, but relies on in-context learning instead of parameter updates. As can be seen from the result, this leads to significantly lower performance than ToSFiT. As FIBO keeps the generated candidate solutions as generation context, it quickly runs out of memory and compute scales quadratically with the number of rounds, forcing us to compare against a top-10 truncated version of FIBO in Figure 4(d).
Exploits Promising Regions Explores using Entropy Explores using Optimism Conditions Model Generation Instruction Following Agnostic
Unguided Generation no no no no yes
Post-Generation TS yes yes yes no yes
Evolutionary Search yes no no yes no
Actor Critic yes no no yes yes
Soft Actor Critic yes yes no yes yes
FIBO yes yes yes yes no
ToSFiT yes yes yes yes yes
Table 1: Among all LLM-based methods considered for optimization, only ToSFiT explores using optimism, conditions model generation for improved efficiency, and does not require the LLM to execute complex procedures via instruction following which sets requirements on the model size.
Scaling to Larger Models
Contrast the top row of experiments in Figure 4 with the bottom row (in Protein Search ProtGPT-2 is used twice for ToSFiT as well as all baselines without instruction following due to the absence of larger variants thereof). While increasing the model size is highly beneficial to FAQ Response Refinement, a natural language task, the tasks of Protein Search and Quantum Circuit Design profit far less. This matches the observations by Romera-Paredes et al. (2024) for code discovery. Indeed, novel discovery often requires leaving the training data manifold, i.e., the generative policy does not have to be tightly concentrated on the most promising solutions.
5.4 ToSFiT Remains Effective in Batched Bayesian Optimization
5.5 ToSFiT Trades Off Computational Efficiency with Sample Efficiency
(a) Latency on fully batched ($B=b=16$) Protein Search. FIBO and Evolutionary Search (LLM) are each tested with Qwen3-0.6B (fast) and Qwen3-8B (slow).
(b) The sample efficiency of ToSFiT on Protein Search is improved by increasing the number of steps of gradient ascent per round (denoted by $c$).
Figure 7: ToSFiT has state-of-the-art computational efficiency (a) and can trade off computational efficiency with sample efficiency (b).
Consider Figure 7(a), which explores the computational efficiency of ToSFiT. We do not plot the Actor Critics and Post-Generation TS, as their computational cost matches, respectively, ToSFiT and Unguided Generation with strictly worse reward. For $10^{\prime}000$ rounds FIBO runs out of memory in its standard implementation, thus we once again implement FIBO attending only to the top-10 candidate solutions and their scores. This also decreases its compute costs drastically at the cost of limiting its sample efficiency. Note that the strong sample efficiency of ToSFiT compensates for the additional compute cost introduced by model fine-tuning, resulting in state-of-the-art computational efficiency. Next, consider Figure 7(b). By investing additional compute through multiple steps of gradient ascent per round, the sample efficiency of ToSFiT can be further improved if needed.
5.6 ToSFiT Benefits from Strong Priors and Requires Careful Fine-Tuning
Lastly, we validate our theoretical insights. Theorem 1 highlights the importance of initializing the sampling policy with strong prior knowledge. Figure 8 provides empirical support: masking the number of qubits in the prompt (ToSFiT weak context) leads to a significantly worse reward than using the full prior context (ToSFiT). Note that uniform initialization fails to produce valid code. The theorem further suggests cautious adaptation to avoid drifting too far from the initialization. As shown in Figure 9, a large learning rate may initially improve performance by closely following VBOS, but eventually leads to forgetting the prior and stagnating as optimization becomes harder.
6 Related Work
As a remedy for the intractable maximization of acquisition functions over combinatorially large discrete domains, Bal et al. (2025) assume a factorization into the cartesian product of much smaller domains. They then identify local maxima as configurations where changing any of the factors individually decreases the acquisition function, and query the black-box function with the highest scoring game equilibrium. In contrast to ToSFiT, the domain must be structured to admit a factorization into cartesian products, and their optimization strategy does not leverage priors from LLMs.
There is extensive prior work on relaxing Bayesian optimization over discrete space to continuous space using variational autoencoders (Kusner et al., 2017; Lu et al., 2018; Griffiths and Hernández-Lobato, 2020; Notin et al., 2021; Lee et al., 2025). Acquisition function maximization is then tractably performed via gradient ascent in the continuous latent space. However, these methods require a task-specific pre-training phase for the autoencoder, whereas ToSFiT can be applied zero-shot through prompt-conditioning, thus benefiting from large-scale industrial LLM pretraining.
Finally, Ranković and Schwaller (2025) learn deep kernel features for Bayesian optimization with Gaussian processes. By treating the neural feature map as a hyperparameter, they use gradient ascent to maximize the marginal likelihood of observations during Bayesian optimization in a fully online fashion. Experimentally, they find much better optimization trajectories compared to fixed feature maps. That said, they do not address the intractability of acquisition function maximization. Their method is complementary to ToSFiT and could be integrated to learn task-specific embeddings.
7 Conclusion
We demonstrate that Thompson sampling can be efficiently scaled to large unstructured discrete domains by parameterizing the probability of maximality with a generative policy. Our theoretical analysis reveals that VBOS already leverages the kernel structure of Gaussian processes and achieves the strong regret bound of standard Thompson sampling and GP-UCB. Extending these results to approximate VBOS highlights the importance of initializing the sampling policy according to strong priors from pre-training and adapting it cautiously to maintain prior knowledge. These insights are supported by empirical results across three diverse tasks. Within a collection of methods covering Bayesian optimization, reinforcement learning, as well as evolutionary search, and to statistical significance, ToSFiT exhibits simultaneously state-of-the-art sample efficiency and computational efficiency. Together, our findings demonstrate the potential of combining foundation models with principled Bayesian optimization to tackle complex, discrete search problems.
7.1 Limitations and Future Work
To ensure controlled comparisons, we evaluate ToSFiT under fixed feature maps—either derived from pre-trained embeddings or manually designed. A promising direction for future work is to learn deep, task-adaptive embeddings jointly with the Gaussian process, as in Ranković and Schwaller (2025), or replace the GP with more expressive reward models such as Bayesian neural networks or ensembles (Lakshminarayanan et al., 2017). Moreover, to reduce the computational and memory overhead introduced by fine-tuning, one could restrict updates to the last few layers of the generator. Finally, alternative strategies for optimizing the VBOS objective $\mathcal{V}(\pi)$—such as in-context conditioning—may offer a lightweight alternative to weight adaptation.
Ethics Statement
Reproducibility Statement
Apart from the code provided in the supplementary materials, pseudocode for ToSFiT is given in Algorithm 1. The baselines are also implemented in the code in a self-contained manner. The experimental environments for the three tasks considered is also provided in the code. Moreover, the provided README.md gives a step-by-step tutorial to reproduce all results in the paper. The hardware and software for running the experiments is described in detail in Section B of the Appendix. Furthermore, Section B.3 details the hyperparameter sweep conducted during development of this paper. In the evaluation, statistical significance is indicated via mean and standard error under $25$ seeds, set to $0,\ldots,24$. All theoretical statements are self-contained and include the complete set of assumptions. Both detailed proofs as well as proof sketches for the main results are provided in Section E of the Appendix. We distinguish between novel claims and prior statements by citing the appropriate works.
Acknowledgment
Appendix
Appendix A Additional Experimental Results
A.1 Computational Cost of ToSFiT
Bayesian optimization is designed to trade the number of evaluations of an expensive reward function against the computationally demanding computation of the reward posterior. In this spirit, we emphasize sample efficiency as the main consideration. Nevertheless, as measured in wall-clock time in Table 2, model fine-tuning only incurs a minor latency overhead (here $19\%$). As a result, despite more cost per round ToSFiT reaches higher reward scores at a fixed computational budget, i.e., it is computationally more efficient than all baselines (see Figure 7(a)).
Autoregressive Generation Model Fine-Tuning GP Update ToSFiT
$1.08\pm 0.05$ s $0.21\pm 0.01$ s $0.33\pm 0.06$ ms $1.29\pm 0.05$ s
Table 2: Wall-clock time per step of fully-batched ToSFiT ($b=B=16$) on an A100 GPU for Protein Search. The majority of the runtime is spent on candidate generation.
Note that in the fully-batched setting considered here all generations are both used to estimate VBOS gradients and to provide observational feedback to the reward model. The 19% overhead observed in Table 2 stems from the additional non-auto-regressive backward pass during model fine-tuning. This computational overhead results in fewer generated candidates at a fixed budget.
A.2 Generation Diversity
A common failure point of reward-based fine-tuning is diversity collapse of the policy. As can be observed in Figure 10, Actor Critic (Barto et al., 1983) is strongly at risk. In contrast, Soft Actor Critic (Haarnoja et al., 2018) retains diversity by directly regularizing the loss with an additional entropy term. Similarly, the objective of ToSFiT contains the additional term $\sqrt{-2\ln\pi_{x}^{\theta}}\cdot\sigma_{x}$, which can be understood as entropy regularization targeted toward underexplored regions where $\sigma_{x}$ is large. As shown in Figure 10, this results in highly robust diversity trajectories. Moreover, in contrast to constant entropy regularization, such an uncertainty-aware exploration-exploitation tradeoff increases diversity in case of large uncertainty, and decreases diversity given sufficient observations.
@@ A.2 Generation Diversity
Figure 10: ToSFiT directs exploration by incorporating uncertainty-aware entropy regularization via $\mathbb{E}_{x\sim\pi^{\theta}}[\sqrt{-2\ln\pi_{x}^{\theta}}\cdot\sigma_{x}]$. It avoids the diversity collapse of Actor Critic and is more stable than Soft Actor Critic (see Subfigure 10(c)). Moreover, by steering generation diversity with the reward uncertainty $\sigma_{x}$, ToSFiT enables both phases of exploration and phases of exploitation. Note that the behavior of each method yields distinct observation histories on top of loss discrepancies.
A.3 Loss and Risk of Overfitting
Figure 11: Evolution of the VBOS objective during Quantum Circuit Design using a Qwen2.5-Coder-1.5B.
The loss function in ToSFiT is set to the negative VBOS objective, i.e., is a lower bound on $\max_{x\in\mathcal{X}}R_{x}$. In contrast to training on a fixed dataset, the loss function evolves during optimization as more data is observed. Thus, the optimization dynamics typically do not lead to a monotonous increase in the objective, see Figure 11. Such non-stationarity leads to incorrect gradient statistics for optimizers like Adam (Kingma and Ba, 2015). Thus, we adopt vanilla stochastic gradient descent. Moreover, in contrast to a fixed dataset, there is no risk of overfitting the policy: the optimal policy is the one that maximizes the VBOS objective, as proven in Theorem 1. Indeed, both entropy regularization and uncertainty in the reward function are directly integrated into the loss. Overfitting of the reward surrogate $\mu_{x}\pm\sigma_{x}$ may still occur, but can be mitigated by adopting ensembles or Gaussian processes.
Appendix B Experimental Details
To ensure reproducibility, we provide the setup for each of the experiments in Section 5. All experiments were conducted on compute nodes with an AMD EPYC 7763 64-Core CPU, 2TB RAM, and an NVIDIA A100-SXM4 GPU (80GB), running on Red Hat Enterprise Linux 9.4 with CUDA 12.4. Our Pytorch-based code (Paszke et al., 2019) is provided in Supplementary Materials, including experimental configuration files and an environment file specifying the necessary Python packages and their versions.
B.1 Baselines
For Unguided Generation, we simply generate and evaluate candidates according to the hyperparameters detailed in Section B.2.
For Post-Generation TS, we first generate a fixed candidate pool using the large language model with the same hyperparameters, where the pool size is set to the number of steps of Bayesian optimization. Then, we conduct standard Thompson sampling on the candidate subset using a Gaussian process surrogate model derived from the (deep) task-dependent embeddings.
For ToSFiT, we generate and evaluate candidates according to the same hyperparameters, but additionally follow Algorithm 1 to perform weight updates for consistency between the policy and the surrogate reward model, the same Gaussian process as in Post-Generation TS.
The Actor Critic baseline follows ToSFiT, but maximizes $\mathbb{E}_{x\sim\pi^{\theta}}[\mu_{x}]$ rather than $\mathcal{V}$.
Soft Actor Critic additionally adds entropy regularization to Actor Critic, i.e., maximizes $\mathbb{E}_{x\sim\pi^{\theta}}[\mu_{x}-\alpha\ln\pi^{\theta}_{x}]$. The hyperparameter $\alpha$ is swept for each experimental setting.
For the Evolutionary Search baselines, we generate new candidates by selecting two parents in each round independently via tournament selection with tournament size $3$ and applying a crossover and mutation operation. We only let the $10$ fittest members of the population survive. The initial population has one entry that acts as an example for the desired format. For the three tasks considered, the initial entry is set to
FAQ: How do I reset my password?
Q: How do I reset my password?
A: To reset your password, follow these steps:
Log in to your account (if you have access to the platform or
service you’re using).
Locate the ’Forgot Password’ or ’Reset Password’ link, usually
found in the login form or menu.
Enter your email address or username associated with your account.
Follow the instructions sent to your email (or the confirmation
screen).
Complete the password reset process by entering a new password
and confirming it.
Submit the form to finalize the reset.
If you’re unable to locate the "Forgot Password" option,
contact the support team for assistance.
MINDLLDISRIISGKMTLDRAEVNLTAIARQVVEEQRQAAEAKSIQLLCSTPDTNHYVFG
DFDRLKQTLWNLLSNAVKFTPSGGTVELELGYNAEGMEVYVKDSGIGIDPAFLPYVFDRF
RQSDAADSRNYGGLGLGLAIVKHLLDLHEGNVSAQSEGFGKGATFTVLLPLKPLKRELAA
VNRHTAVQQSAPLNDNLAGMKILIVEDRPDTNEMVSYILEEAGAIVETAESGAAALTSLK
SYSPDLVLSDIGMPMMDGYEMIEYIREWKTTKGG
qc.cx(0, 1)
Evolutionary Search (Character) applies crossover by cutting each parent in half at a uniformly chosen location and stitching the pieces together. Mutation is performed through random substitution (at rate $0.05$ per position) of the characters of the candidate solution with a letter from the vocabulary ACDEFGHIKLMNOPQRSTUVWY for Protein Search and one of the 95 printable ASCII characters for FAQ Response Refinement and Quantum Circuit Design. After substitution, random insertion & deletion of characters (at rate $0.01$) is performed.
Evolutionary Search (LLM) performs crossover and mutation in context through instruction following of the subsequent system prompt
You are conducting evolutionary search in context. You are
provided two candidate solutions. Propose a new distinct solution
by combining the candidate solutions and mutating the result.
Your novel candidate solution must be enclosed by
<candidate> </candidate>. Never repeat previous solutions. Your
search space is over FAQ responses to the question "How do I reset
my password?". /no_think
[...] Your search space is over amino acid sequences, with each
amino acid represented as a single-letter code./no_think
[...] Your search space is over 6-qubit Qiskit quantum circuits,
already prepared with ‘qc = QuantumCircuit(6)‘.
Finally, FIBO implements Thompson sampling fully in context by attending to a list of evaluated candidate solutions, initially set to the same examples provided to Evolutionary Search (LLM). The algorithm implements Thompson sampling by adopting the following system prompt:
You are conducting Bayesian optimization (Thompson sampling) fully
in context. You are provided a list of candidate solutions and the
rewards achieved by these solutions. Propose a new distinct
solution that maximizes the reward. Your novel candidate solution
must be enclosed by <candidate> </candidate>. Never repeat
previous solutions. Your search space is over FAQ responses to the
question "How do I reset my password?". /no_think
[...] Your search space is over amino acid sequences, with each
amino acid represented as a single-letter code./no_think
[...] Your search space is over 6-qubit Qiskit quantum circuits,
already prepared with ‘qc = QuantumCircuit(6)‘.
B.2 Experimental Details Shared Across Experiments
@@
ToSFiT requires minimal task-specific hyperparameter tuning. Across all experiments with a Gaussian process reward model, a burn-in period of $m=16$ steps is adopted and an exploration bonus of $4.0$ is applied to the prior amplitude inferred via MLM. The noise-to-amplitude ratio $\sigma_{nar}$ is always set to $0.01$, reflecting the absence of noise in the measurements, but ensuring stable inversion of the observation covariance matrix. Unless stated otherwise, the number of steps of gradient ascent per step of Bayesian optimization, $c$, is fixed at $1$ to maximize computational efficiency. To ensure stable estimation of VBOS gradients, the generation batch size $B$ is always set to $16$ with temperature set to $1.0$. We conduct unbatched Bayesian optimization ($b=1$) unless otherwise specified. Learning rates are the only hyperparameters tuned per task: $1.0\times 10^{-7}$ for FAQ refinement, $1.0\times 10^{-5}$ for Protein Search, and $5.0\times 10^{-6}$ for Quantum Circuit Design. The baseline Soft Actor Critic requires an additional entropy regularization coefficient as hyperparameter and is swept across $\{0.01,0.1,1.0,10.0\}$ for each experiment. We report on the best configuration.
@@
To reduce the memory overhead of ToSFiT compared to Unguided Generation (and in light of Section A.3), we do not adopt parameter-wise optimization statistics (Kingma and Ba, 2015). Instead, we use vanilla stochastic gradient descent with a fixed learning rate and without momentum. We do not store the gradients, but instead apply them directly during the backward pass. Also, while we perform fully-batched auto-regressive generation, the forward & backward passes to compute $\tfrac{d}{d\theta}\ln\pi_{\theta}(x_{i})$ are performed in unbatched fashion with checkpointing. This crucially brings down memory consumption with little impact on runtime. All experiments were run on a single NVIDIA A100 GPU with 80 Gigabytes of VRAM.
B.3 Choice of Hyperparameters
As mentioned above, ToSFiT requires minimal task-specific hyperparameter tuning. Indeed, only the learning rate is tuned for each task, based on one-dimensional grid search with search locations $\{5.0\times 10^{-5},1.0\times 10^{-5},5.0\times 10^{-6},1.0\times 10^{-6},5.0\times 10^{-7},2.0\times 10^{-7},1.0\times 10^{-7},5.0\times 10^{-8}\}$. Burn-in periods of $m\in\{4,8,16\}$ were tried out with $16$ being sufficiently large to ensure stable marginal likelihood maximization, yet small enough to not significantly impact sample efficiency due to delayed optimization. The exploration bonus was sweeped across the set $\{1.0,2.0,4.0,8.0,16.0\}$ on a toy task. A bonus of $4.0$ provided the best trade-off between exploration and exploitation, and generalized to strong performance on the reported experiments. The number of steps of gradient ascent per step of Bayesian optimization, $c$, was kept to $1$ in order to minimize the computational overhead compared to Unguided Generation. For the generation batch size $B$, the values $\{4,8,16,32,64\}$ were tried out. $16$ was the smallest size where GPU utilization became acceptably large during auto-regressive generation. However, empirically $B$ did not significantly impact the simple regret trajectories during Bayesian optimization, suggesting that it could possibly be lowered for alternative models and/or hardware. The baseline Soft Actor Critic requires an additional entropy regularization coefficient as hyperparameter and is swept across $\{0.01,0.1,1.0,10.0\}$ for each experimental setup.
B.4 FAQ Refinement